In rectangle ABCD, side AB=6, side BC=9, and diagonal AC is drawn. Find angle CAD to the nearest degree.

Answer by scott8148(6628)

CAD is congruent to ACB

tan(ACB)=AB/BC __ tan(ACB)=6/9 __ tan(ACB)=2/3